### v1

aHigh Energy Physics Department, Institute for Mathematics, Astrophysics,
and Particle Physics, Radboud University, Nijmegen, The Netherlands
bSchool of Fundamental Physics and Mathematical Sciences,
Hangzhou Institute for Advanced Study, UCAS, Hangzhou 310024, China
cInstitute of Theoretical Physics, Chinese Academy of Sciences, Beijing 100190, China
dUniversity of Chinese Academy of Sciences, Beijing 100049, China
eInternational Centre for Theoretical Physics Asia-Pacific, Beijing/Hangzhou, China
The requirements of unitarity and causality lead to significant constraints on the Wilson coefficients of an EFT expansion, known as positivity bounds. Their standard derivation relies on the crucial assumption of polynomial boundedness on the growth of scattering amplitudes in the complex energy plane, which is a property satisfied by local QFTs, and by weakly coupled string theory in the Regge regime. The scope of this work is to clarify the role of locality by deriving generalized positivity bounds under the assumption of exponential boundedness, typical of non-local QFTs where the Froissart-Martin bound is usually not satisfied. Using appropriately modified dispersion relations, we derive new constraints and find regions in the EFT parameter space that do not admit a local UV completion. Furthermore, we show that there exist EFTs that satisfy IR causality and at the same time can admit a non-local UV completion, provided that the energy scale of non-locality is of the same order or smaller than the EFT cutoff. Finally, we provide an explicit example of a non-perturbative amplitude that simultaneously satisfies the properties of exponential boundedness, unitarity and causality.
$\,{}^{\star}$ luca.buoninfante@ru.nl
$\,{}^{\dagger}$ shaolongqi22@mails.ucas.ac.cn
$\,{}^{\ddagger}$ tokareva@ucas.ac.cn
1 Introduction
Scattering amplitudes are very powerful tools for connecting the ultraviolet (UV) aspects of a fundamental theory with its properties and observables at low energies. To constrain the parameter space of viable effective field theories (EFTs) it is often not necessary to know the precise form of the UV completion. In many cases it is sufficient to assume certain physical properties in the UV regime, which can be translated into mathematical requirements on the $S$-matrix, from which constraints on low-energy physics can then be derived [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] (see also  [54] for a recent review). For example, standard requirements that are usually assumed are those of Lorentz invariance, unitarity and causality. These properties are encoded in a set of precise mathematical conditions, such as positivity of the imaginary part of the partial waves and analyticity of the $2\rightarrow 2$ scattering amplitude in the complex energy plane [55, 56, 57, 58].
A further assumption that is typically made on the UV completion is that of locality, which states that the growth of the scattering amplitude is polynomially bounded in the complex energy plane. For local gapped theories it can be proven that the amplitude for fixed physical momentum transfer is polynomially bounded by $|s|^{2}$ in the whole complex plane; this is known as Froissart-Martin bound [^1] It is worth to mention that Froissart [59] and Martin [60] initially showed that the amplitude was bounded by $|s|$ in the forward limit $t=0.$ Subsequently, Jin and Martin [61] showed that the amplitude for generic physical momentum transfer $t\leq 0$ is bounded by $|s|^{2}.$ In any case, in this work we refer to the latter as Froissart-Martin bound. [59, 60, 61]. Although for theories with massless states no rigorous general proof of this bound exists, it is common to assume that the growth of the amplitude for fixed physical $t$ is bounded by $|s|^{2}$ in the whole complex plane [^2] This can actually be verified for graviton-mediated amplitudes in $d>4$, using the method of eikonal resummation [62, 63].:
$|A(s,t\leq 0)|<B|s|^{2}\,,\qquad|s|\rightarrow\infty\,,$ (1)
where $B$ is some subdominant factor that can also depend on $s$.
To clarify the role of locality, it is instructive to investigate what happens when we allow for a non-polynomial growth of the scattering amplitude, i.e., for violations of the Froissart-Martin bound. In particular, we are going to consider amplitudes that are exponentially bounded in the complex $s$ plane for fixed physical momentum transfer:
$|A(s,t\leq 0)|<C|s|^{2}e^{\alpha|s|^{2g}}\,,\qquad|s|\rightarrow\infty\,,$ (2)
where $\alpha$ and $g$ are two constant positive parameters, and $C$ a subdominant factor that can also depend on $s$.
QFTs in which scattering amplitudes are characterized by this type of exponential behavior could be related to Jaffe’s classification [64, 65, 66, 67, 68]: if $g<1/4$ the corresponding QFT is said to be localizable, whereas if $g\geq 1/4$ the QFT is said to be non-localizable. In particular, when $g=0$ we recover the standard local QFT description with polynomially bounded amplitudes. The name non-localizability is related to the fact that the fields must be delocalized through smeared test functions in order to make the position-space correlators (such as the Wightman function) finite and well-defined. In what follows we use the words local (non-local) and localizable (non-localizable) interchangeably.
The aim of this paper is to understand what type of constraints can be derived on the Wilson coefficients of the low-energy EFT expansion if the UV completion is assumed to be non-localizable according to (2). As a case study, we consider a gapless scalar theory whose full $2\rightarrow 2$ scattering amplitude $A(s,t)$ vanishes at the point $(s,t)=(0,0)$ and that in the forward limit is bounded as $A(s,t=0)<s^{2}e^{-\alpha s^{2g}}$ in the entire complex plane. Implementing modified dispersion relations that are suitable for exponentially bounded amplitudes, we will show that the standard positivity bounds [3, 54] are generalized and become dependent on the parameters $\alpha$ and $g$ in Eq. (2). In particular, we focus on the case $g=1.$ We will find that the generalized positivity bounds not always overlap with the standard ones but, depending on the value of $\alpha,$ there can exist also regions in the EFT parameter space that do not admit a local UV completion. We further compare these positivity constraints with those arising from the low-energy causality requirement and find large regions in the EFT parameter space where the theory is both causal and unitary, but can be non-localizable. After presenting a general investigation of positivity and causality bounds, we will construct explicit examples of non-perturbative scattering amplitudes that satisfy analyticity, crossing symmetry and partial-wave unitarity, but are not polynomially bounded.
The work is organized as follows.
Sec. 2:
We introduce the method of dispersion relations suitably adapted to the case of amplitudes that are exponentially bounded in the complex $s$ plane. In particular, we consider the case of forward limit, i.e. zero-momentum transfer, and specialize the study to $g=1$.
Sec. 3:
We derive the generalized positivity bounds in the forward limit by constraining the coupling constants $g_{4},$ $g_{6}$ and $g_{8}$ in the EFT expansion. We determine the dependence of the result on the value of the non-local scale $\alpha$, and make a comparison with the local case $\alpha=0.$
Sec. 4:
We generalize the analysis to the case of non-zero momentum transfer and derive further generalized positivity bounds that also constrain the coupling $g_{3}$ in the EFT expansion. In this case, additional assumptions – such as crossing symmetry – are necessary in order to write down a well-defined dispersion relation.
Sec. 5:
We compare the regions in the space of EFT parameters constrained by the positivity bounds with the regions limited by the requirement of low-energy causality. We show that non-local UV completions can be compatible with the requirement of causality only if the energy scale of non-locality is of the same order or smaller than the EFT cutoff.
Sec. 6:
We comment on the implications of our results. In particular, we remark that our findings challenge the usually-claimed uniqueness of string theory as the only UV completion of gravity that is consistent with the requirements of unitarity and causality. Finally, we draw our conclusions.
App. A:
In this appendix we construct an explicit example of a non-perturbative amplitude that satisfies the properties of analyticity, causality, partial-wave unitarity, but violate polynomial boundedness.
App. B:
More mathematical details on the analytic derivation of the bound for the coupling $g_{3}$ are provided.
Conventions and notations.
Throughout the entire paper, we work with the mostly plus metric signature ($+$)(-$+$)( - + + + ) and adopt the Natural units system $(\hbar=1=c).$ The central object of our investigation is the elastic $2\rightarrow 2$ scattering amplitude $A(s,t,u),$ where $s=-(p_{1}+p_{2})^{2},$ $t=-(p_{1}-p_{3})^{2},$ $u=-(p_{1}-p_{4})^{2}$ are the Mandelstam variables, $p_{1},$ $p_{2}$ are the ingoing momenta and $p_{3},$ $p_{4}$ are the outgoing momenta. In particular, $s$ is equal to the centre-of-mass energy squared, while $t$ is the momentum transfer squared. Since in this work we only focus on a gapless theory, we have $s+t+u=0.$ The scattering amplitude can be chosen to be a function of two Mandelstam variables only, e.g. $s$ and $t$, thus $A(s,t,u)=A(s,t)$.
2 Dispersion relations for exponentially bounded amplitudes
The central object of our discussion is the $2\rightarrow 2$ scattering amplitude $A(s,t)$ in some gapless scalar theory. In particular, we assume that its low-energy behavior can be captured by the following EFT expansion:
$A_{\rm EFT}(s,t)=$ $\,g_{2}(s^{2}+t^{2}+u^{2})+g_{3}stu+g_{4}(s^{2}+t^{2}+u^{2})^{2}$ (3)
$+g_{5}(stu)(s^{2}+t^{2}+u^{2})+g_{6}(s^{2}+t^{2}+u^{2})^{3}+g^{\prime}_{6}(stu)^{2}+\cdots\,,$
which is manifestly crossing symmetric and vanishes at $(s,t)=(0,0)$; the couplings $g_{n}$ are known as Wilson coefficients.
Polynomially bounded amplitude.
In the past literature [3, 54], the Wilson coefficients $g_{n}$ have been constrained assuming that the UV completion is local, i.e. the amplitude is polynomially bounded in all complex directions in the $s$ and $t$ planes. This requirement together with partial-wave unitarity implies the Froissart-Martin bound in Eq. (1). In addition to these ingredients, the assumption of analyticity of the amplitude allows to derive various dispersion relations at fixed $t\leq 0$ connecting the Wilson coefficients in $A_{\rm EFT}$ with positive contributions coming from the discontinuities along the branch cuts of the amplitude. If we choose the contour $\gamma_{1}\cup\gamma_{2}$ in Fig. 1, we can write [6, 24]
$\int_{\gamma_{1}\cup\gamma_{2}}\frac{{\rm d}\mu}{2\pi i}\frac{A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}=0\,,$ (4)
where $s_{1}$ and $s_{2}$ are two subtraction points. Subsequently, by deforming the integration contour as shown in Fig. 1, Eq. (4) can be recast as
$\int_{C_{1}\cup C_{2}}\frac{{\rm d}\mu}{2\pi i}\frac{A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}=$ $\,(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}$ (5)
$+(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-u_{1})(\mu-u_{2})(\mu-u)}\,,$
where
${\rm Disc}_{s}A(\mu,t)=\lim_{\epsilon\rightarrow 0^{+}}\frac{1}{2i}\left[A(\mu+i\epsilon,t)-A(\mu-i\epsilon,t)\right],$ (6)
and we have used crossing symmetry to write ${\rm Disc}_{s}A(-\mu-t,t)=-{\rm Disc}_{s}A(\mu,t)$ and defined analogue subtraction points in terms of the Mandelstam variable $u,$ i.e. $u_{1}=-s_{1}-t$ and $u_{2}=-s_{2}-t.$ The scale $\Lambda$ represents the cutoff scale of the EFT.
Figure 1: Integration contour in the complex $\mu$ plane used to derive the twice-subtracted dispersion relations in eqs. (7), (9) and (11). The initial contour $\gamma_{1}\cup\gamma_{2}$ on the left is taken of infinite radius. It is then deformed into the contour on the right, where the circle $C_{1}\cup C_{2}$ is centered at $\mu=-t/2.$ The boxes denote the two branch points of the $s$-channel (from $\mu=0$ to +$\infty$+ ∞) and $u$-channel (from $\mu=-t$ to $\infty$- ∞) branch cuts, respectively. The point $s$, $s_{1}$ and $s_{2}$ are the poles for which the residue theorem is used to derive the dispersion relations.
By computing the integral on the left-hand-side of Eq. (5) through the residue theorem, we obtain the twice-subtracted dispersion relation
$A_{\rm EFT}(s,t)=$ $\,(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}$ (7)
$+(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-u_{1})(\mu-u_{2})(\mu-u)}+a(t)+sb(t)\,,$
where
$a(t)\equiv\frac{1}{s_{1}-s_{2}}\left[s_{1}A_{\rm EFT}(s_{2},t)-s_{2}A_{\rm EFT}(s_{1},t)\right]\,,\qquad b(t)\equiv\frac{1}{s_{1}-s_{2}}\left[A_{\rm EFT}(s_{1},t)-A_{\rm EFT}(s_{2},t)\right]$ (8)
are the subtraction terms. Note that we wrote $A_{\rm EFT}(s,t)$ because the left-hand-side of Eq. (7) satisfies the equation only for $s<\Lambda^{2}.$ This aspect is what allows us to make the connection with the low-energy EFT expansion (3). In fact, more precisely, the left-hand-side of Eq. (7) reads $A(s,t)|_{s<\Lambda^{2}}\simeq A_{\rm EFT}(s,t).$
To work with simpler formulas, in what follows we choose that the two subtraction points coincide and are equal to zero, that is, $s_{1}=s_{2}=0.$ In this way, the dispersion relation (7) becomes
$A_{\rm EFT}(s,t)=s^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{\mu^{2}(\mu-s)}+s^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu+t)^{2}(\mu-u)}+a(t)+sb(t)\,,$ (9)
where now the subtraction terms read
$a(t)=A_{\rm EFT}(0,t)\,,\qquad b(t)=\left.\frac{\partial A_{\rm EFT}(s,t)}{\partial s}\right|_{s=0}\,.$ (10)
Taking the second and higher order derivatives at $s=0$ of the dispersion relation (7) or (9), and combining the expressions into a set of integral inequalities, one can derive constraints on $g_{n}$ [54].
Exponentially bounded amplitude.
In this work, we are interested in non-local UV completions in which the amplitude does not satisfy the property of polynomial boundedness. In particular, for fixed momentum transfer we assume that its growth for $|s|\rightarrow\infty$ is bounded as in Eq. (2). In this case the requirements of analyticity and partial-wave unitarity can still be satisfied, that is, the only pole singularities and branch cuts are confined to the real axis and the discontinuity across the branch cut is positive. However, in this case the integral over $\gamma_{1}\cup\gamma_{2}$ in Eq. (4) could blow up. [^3] It is worth to mention that for strictly localizable QFTs, in which the amplitude is bounded as (7) with $g<1/4,$ it can actually be shown that the amplitude cannot grow faster than a polynomial in the complex $s$ [67]. In this paper, instead, we are interested in non-localizable QFTs with $g>1/4$ (in particular $g=1$), for which a similar result does not necessarily apply. Therefore, modified dispersion relations that take into account a non-localizable exponential growth are needed. This means that we need to implement a modified dispersion relation. Let us start considering the forward limit $t=0,$ and later on we will address the case of non-vanishing momentum transfer that will require a more careful treatment.
When $t=0$, we can assume that the amplitude is bounded as $A(s,t=0)<s^{2}\exp{(-\alpha s^{2g})}$ for any $s\in\mathbb{C},$ up to subdominant factors. This means that along the real axis the amplitude is exponentially suppressed in the forward limit, and can grow exponentially along some direction in the complex $s$ plane. To write down a dispersion relation for this type of amplitude, we can adopt the following strategy. We multiply the amplitude by $\exp{(\alpha s^{2g})}$ and define $A^{\prime}(s,t)\equiv\exp{(\alpha s^{2g})}A(s,t)$, which will be bounded by $|s|^{2}$ in the whole complex $s$ plane. In other words, we compensate the essential singularity at infinity, so that we can write down dispersion relations for $A^{\prime}(s,t=0)$ that are analogue to Eq. (9). Indeed, we have
$\exp{(\alpha s^{2g})}A(s,0)=\,s^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{\rm{Disc}_{s}A^{\prime}(\mu,0)}{\mu^{2}(\mu-s)}+s^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{\rm{Disc}_{s}A^{\prime}(\mu,0)}{\mu^{2}(\mu+s)}+a^{\prime}(0)+sb^{\prime}(0)\,,$ (11)
where the subtraction terms $a^{\prime}(0)$ and $b^{\prime}(0)$ have the same form as those in eq. (10), but with $A$ replaced by $A^{\prime}$. However, since our amplitude satisfies $A_{\rm EFT}(0,0)=0$ and $\partial_{s}A_{\rm EFT}(s,0)|_{s=0}=0,$ the two subtraction terms actually vanish in the forward limit, i.e. $a^{\prime}(0)=0=b^{\prime}(0).$
The relation between the Wilson coefficients in $A_{\rm EFT}$ and the positive moments – i.e. the derivatives with respect to $s$ evaluated at $s=0$ of the amplitude – will now be modified due to additional $\alpha$-dependent terms. In the next section we show how this affects the resulting EFT constraints.
We will also provide explicit examples of amplitudes that decay exponentially in the UV along physical directions and, at the same time, satisfy partial-wave unitarity, causality and crossing symmetry. For instance, the expression
$\begin{split}A(s,t,u)\sim\,C&\left[e^{-\alpha s^{2}}\log{(\sqrt{\alpha}s)}\log{(-\sqrt{\alpha}s)}+e^{-\alpha t^{2}}\log{(\sqrt{\alpha}t)}\log{(-\sqrt{\alpha}t)}+\right.\\ &\left.+e^{-\alpha u^{2}}\log{(\sqrt{\alpha}u)}\log{(-\sqrt{\alpha}u)}\right]\end{split}$ (12)
represents a possible UV behavior that is consistent with partial-wave unitarity for $s\gtrsim 1/\sqrt{\alpha}$ and that could be matched to the logarithmic running at low energies. In App. A we examine further examples, including amplitudes that are finite at the point $(s,t)=(0,0),$ which are therefore consistent with our working assumptions.
3 Positivity bounds at $t=0$
Let us now derive the positive bounds in the forward limit. Hereafter we study the non-localizable case with $g=1$. Taking two derivatives of the dispersion relation (11) with respect to $s$ and evaluating them at $s=0$, we obtain
$g_{2}=\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\rm{Disc}_{s}A^{\prime}(\mu,0)\frac{1}{\mu^{3}}\,.$ (13)
Unitarity implies the positivity of the last integral, thus we obtain the first constraint on the lowest Wilson coefficient, i.e. $g_{2}>0$, which remains the same as in the standard local case.
If we define
$\frac{1}{\mu^{m}}\right>\equiv\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}{\rm Disc}_{s}A^{\prime}(\mu,0)\frac{1}{\mu^{m}}\,,$ (14)
we can write
[…]
and analogue relations can be obtained for higher-order Wilson coefficients.
Figure 2: (a) Allowed regions in the $\tilde{g}_{4}$-$\tilde{g}_{6}$ plane when $\tilde{\alpha}=0,0.2,0.41,1$. (b) Allowed regions in the three-dimensional parameter space spanned by $\tilde{g}_{4}$, $\tilde{g}_{6}$ and $\tilde{g}_{8}$ when $\tilde{\alpha}=0,0.5,1$.
Working directly with the last relations is not convenient due to the presence of negative contributions on the right-hand-sides. However, we can instead manipulate those equations and work with the following positive combinations of Wilson coefficients,
[…]
$\frac{1}{\mu^{9}}\right>=\,g_{8}+4\alpha g_{6}+\alpha^{2}g_{4}+\frac{\alpha^{3}}{6}g_{2}\,.$
[…]
$0<2\tilde{g}_{4}+\tilde{\alpha}<1,\qquad 0<4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\frac{1}{2}\tilde{\alpha}^{2}<2\tilde{g}_{4}+\tilde{\alpha},\qquad 4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\frac{1}{2}\tilde{\alpha}^{2}\geq(2\tilde{g}_{4}+\tilde{\alpha})^{2},$ (17)
where we have defined the dimensionless quantities $\tilde{\alpha}\equiv\alpha\Lambda^{4},$ $\tilde{g}_{4}\equiv g_{4}\Lambda^{4}/g_{2}$, $\tilde{g}_{6}\equiv g_{6}\Lambda^{8}/g_{2}$, and so on. In the $\alpha\rightarrow 0$ limit, these inequalities reduce to the well-known local case [3]. We also verified that using two Cauchy-Schwartz inequalities to derive the bounds on $\tilde{g}_{4}$ and $\tilde{g}_{6}$ is already sufficient to obtain the most optimal constraints, which is in agreement with other analyses performed in local cases [69].
The regions in the $\tilde{g}_{4}$-$\tilde{g}_{6}$ plane that are allowed by the constraints in (17) are shown in Fig. 2a for different values of $\tilde{\alpha}$. We can track how the allowed region changes as a function of $\tilde{\alpha}$, going from $\tilde{\alpha}=0$ to $\tilde{\alpha}=1.0$. In particular, we found that the regions corresponding to $\tilde{\alpha}\geq\sqrt{2}-1\approx 0.41$ have no overlap with the local case $\tilde{\alpha}=0$.
In Fig. 2b we plotted the allowed regions in the three-dimensional parameter space spanned by $\tilde{g}_{4}$, $\tilde{g}_{6}$ and $\tilde{g}_{8}$.
4 Bounds on $g_{3}$ coupling
Although the above derivation of positivity bounds in the forward limit is quite straightforward, it does not give us any information about the couplings in the EFT amplitude (3) that disappear when $t=0$. Constructing the dispersion relation for $t<0$ requires more details about the structure of the UV amplitude. For example, at finite $t$ it may be not possible to compensate the essential singularities in dispersion relation in a crossing-symmetric way. To avoid these potential obstacles, we assume a crossing-symmetric form of the amplitude which is consistent with partial-wave unitary and is inspired by the explicit examples provided in App. A. In particular, if the amplitude can be represented as a sum of $s$, $t$, and $u$ channels weighted with the corresponding exponentials [^4] Here we assume $F(s,t)$ to be bounded by $s^{2}$ in the entire complex plane, so that it can be expressed in terms of a twice-subtracted dispersion relation.,
$A(s,t,u)=F(t,u)e^{-\alpha s^{2g}}+F(s,u)e^{-\alpha t^{2g}}+F(s,t)e^{-\alpha u^{2g}},$ (18)
we can derive its $s\leftrightarrow u$ crossing-symmetric representation (specializing to $g=1$),
$A_{\rm{EFT}}(s,t)=e^{-\alpha s^{2}}s^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,t)}{\mu^{2}(\mu-s)}+e^{-\alpha u^{2}}u^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,t)}{\mu^{2}(\mu-u)}+a(t)\,,$ (19)
where $A^{\prime}(\mu,t)=A(\mu,t)e^{\alpha\mu^{2}}$. Note that there is no subtraction term $b(t)$, otherwise the $s\leftrightarrow u$ crossing symmetry would be broken.
Moreover, if we further assume full crossing symmetry, we can derive the following expression for $a(t)$:
$a(t)=a(0)+e^{-\alpha t^{2}}t^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,0)}{\mu^{2}}\left(\frac{1}{\mu-t}+\frac{1}{\mu+t}\right)-e^{-\alpha t^{2}}t^{2}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A^{\prime}(\mu,t)}{\mu^{2}(\mu+t)}\,,$ (20)
where $a(0)=0$ in our case since we have $A_{\rm EFT}(0,0)=0.$ We can easily verify that in the limit $t\rightarrow 0$ the dispersion relation (19) consistently recovers (11) which was derived in the forward limit.
Equipped with a dispersion relation that is valid for a physical non-zero $t,$ we can derive constraints on additional Wilson coefficients, e.g. on the coupling $\tilde{g}_{3}$. To do so, we need to implement the partial-wave decomposition for the discontinuities in $d=4$ spacetime dimensions, i.e.
$A(s,t)=\sum_{l\,\rm{even}}n_{l}f_{l}(s)\mathcal{P}_{l}\left(1+\frac{2t}{s}\right)\,,\qquad n_{l}=16\pi(2l+1)\,,$ (21)
from which we get
${\rm Disc}_{s}A(s,t)=\sum_{l\,\rm{even}}n_{l}\rho_{l}(s)\mathcal{P}_{l}\left(1+\frac{2t}{s}\right)\,,$ (22)
where $f_{l}(s)$ is the partial-wave amplitude, $\rho_{l}(s)={\rm Im}f_{l}(s)$ is the spectral density, $\mathcal{P}_{l}(x)$ are the Legendre polynomials, and the angular momentum $l$ only takes even values because we are working with a scalar field theory. If we define
$\left[\frac{1}{\mu}\right]\equiv\sum_{l\,\rm{even}}n_{l}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\rho_{l}(\mu)e^{\alpha\mu^{2}}\frac{1}{\mu},$ (23)
the dispersion relation (19) can be written in the compact form,
$A(s,t)=e^{-\alpha s^{2}}s^{2}\left[\frac{\mathcal{P}_{l}(1+2t/\mu)}{\mu^{2}(\mu-s)}\right]+e^{-\alpha u^{2}}u^{2}\left[\frac{\mathcal{P}_{l}(1+2t/\mu)}{\mu^{2}(\mu-u)}\right]+a(t).$ (24)
Taking derivatives with respect to $s$ and $t$ and setting $s\rightarrow 0$, $t\rightarrow 0$ afterwards, we obtain
$g_{2}=\left[\frac{1}{\mu^{3}}\right],\quad g_{3}=\left[\frac{3}{\mu^{4}}\right]-\left[\frac{2L^{2}}{\mu^{4}}\right],\quad g_{4}=\left[\frac{1}{2\mu^{5}}\right]-\left[\frac{\alpha}{2\mu^{3}}\right].$ (25)
where $L^{2}\equiv l(l+1)$. Crossing symmetry implies that the relation for the coupling $g_{4}$ in (3) can be derived in two different ways. The first way is differentiating four times with respect to $s$ and then taking $s,t\rightarrow 0$. The second way is differentiating twice with respect to $s$ and twice with respect to $t$ and then taking $s,t\rightarrow 0$. Since the two procedures must be equivalent, we get the so-called null constraint reflecting the crossing symmetry of low-energy EFT amplitude,
$\left[\frac{L^{2}(L^{2}-8)}{\mu^{5}}\right]=0\,.$ (26)
Let us make two non-trivial observations about the comparison between the method we are using to derive constraints on $g_{3}$ in non-local theories and the procedure in the case of local theories.
  • The null constraint (26) has the same form as in local theories, except for the different definition of the brackets that are adapted to the case of exponentially bounded amplitudes.
  • The relation (25) between $g_{2}$, $g_{3}$, and $g_{4}$ gets modified only through the one for $g_{4}$ (modulo brackets definition), such that the bound in the non-local case can be obtained from the local one upon the substitution $\tilde{g}_{4}^{\rm local}\rightarrow\tilde{g}_{4}+\tilde{\alpha}/2$.
The state-of-the-art technique to derive the bound on $\tilde{g}_{3}$ is based on the theory of moments and is described in [69, 70]. We briefly review the formalism in App. B and also show that it does not depend on the particular definition of the brackets. In other words, up to additional $\alpha$-dependent terms in some of the relevant relations, the constraints in the $g_{3}$-$g_{4}$ plane in our non-local model can be derived by following the same steps as in the local case [69, 70]. The result is given by the following two sets of inequalities:
I: $\left\{\begin{array}[]{l}\displaystyle-\sqrt{\frac{427}{3}(2\tilde{g}_{4}+\tilde{\alpha})}<\tilde{g}_{3}<3\sqrt{2\tilde{g}_{4}+\tilde{\alpha}}\,,\\[4.2679pt] \displaystyle-\frac{\tilde{\alpha}}{2}\leq\tilde{g}_{4}\leq\frac{961}{2562}-\frac{\tilde{\alpha}}{2}\,,\end{array}\right.$ (29)
II: $\left\{\begin{array}[]{l}\displaystyle-\frac{62}{6}(2\tilde{g}_{4}+\tilde{\alpha})-\sqrt{\frac{320}{9}\left[(2\tilde{g}_{4}+\tilde{\alpha})-(2\tilde{g}_{4}+\tilde{\alpha})^{2}\right]}<\tilde{g}_{3}<3\sqrt{2\tilde{g}_{4}+\tilde{\alpha}}\,,\\[4.2679pt] \displaystyle\frac{961}{2562}-\frac{\tilde{\alpha}}{2}\leq\tilde{g}_{4}\leq\frac{1-\tilde{\alpha}}{2}\,.\end{array}\right.$ (32)
The most optimal bound on $g_{3}$ requires the use of the next order null constraint [69], which will certainly be modified in the non-local case. It is worth mentioning that adding higher-order null constraints would only lead to a very small improvement of the bound in its upper left corner. This does not affect any of our further conclusions and comparisons with causality bounds in the next section.
In Fig. 3a we have shown the allowed regions in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane given by the bounds (29), for different values of $\tilde{\alpha}$. The local constraint is just shifted down allowing for negative values of $g_{4}$, which are forbidden in the standard local case.
Figure 3: (a) Allowed region in the $\tilde{g}_{3}$-$\tilde{g}_{4}$ plane for $\tilde{\alpha}=0,0.5,1$ from (29) and (32). (b) Allowed regions by IR causality bounds (blue region) [71] and bounds on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ in both local ($\tilde{\alpha}=0$) and non-local ($\tilde{\alpha}=1.02$) cases. For $\tilde{\alpha}<1.02$ there is an overlap between EFT constraints and the requirement of IR causality, while larger values of $\tilde{\alpha}$ are in contrast with IR causality.
5 IR causality bounds
The requirement of causality for the low-energy theory imposes a set of bounds on the Wilson coefficients of the EFT expansion (3), that must hold independently of the type of UV completion. Causality can be probed by considering a process in which a plane wave propagates on a classical background created by some local source [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. The causality condition then requires that the presence of the background should not induce superluminal propagation of the plane wave.
However, the uncertainty principle prevents us from detecting very small time advances in the signal propagation, i.e. $|\Delta t|\lesssim 1/\omega,$ where $\omega$ is the energy of the propagating wave. Therefore, unobservable causality violations are still physically allowed because, thanks to their quantum stochastic nature, they do not sum up into macroscopic time advances in signal propagation [86, 87, 71, 88]. This means that we can physically admit sufficiently short time advances that are within the uncertainty principle, i.e. time delay $\Delta t$ can also get very small negative values:
$\Delta t>-1/\omega.$ (33)
[…]
A comprehensive study of causality bounds in the case of shift-symmetric scalar theories based on the condition in Eq. (33) was performed in [71]. We plotted the causality constraints on $\tilde{g}_{3}$ and $\tilde{g}_{4}$ derived in [71] together with our bounds in Fig. 3b. [^6] Authors are indebted to Marianna Carillo-Gonzales for sharing the numerical results of [71] on IR causality bounds. Let us make some important remarks on our results.
  • IR causality bounds are non-compact and are significantly weaker than local positivity constraints on the Wilson coefficients, thus leaving a place in the space of EFT parameters for causal EFTs whose UV completion can only be non-local.
  • As already noticed in the previous section, exponentially bounded amplitudes allow for negative values of $\tilde{g}_{4}$, which are not compatible with local UV completions. On the other hand, Fig. 3b shows that IR causality bounds (which are independent of the type of UV completion) allow for a very small negative $\tilde{g}_{4}$.
  • Furthermore, if the non-local scale $\alpha$ is too large, then the allowed regions derived from the non-local positivity bounds will become incompatible with the IR causality constraints. This way, we get a very interesting constraint on the UV theory imposed by IR requirements:
    $\tilde{\alpha}=\alpha\Lambda^{4}<1.02\quad\Leftrightarrow\quad M\leq(1.02)^{1/4}\Lambda\simeq\Lambda\,,$ (34)
    where the second inequality is written in terms of the non-local energy scale $M\equiv 1/\alpha^{1/4}.$ This means that a causal EFT with cutoff $\Lambda$ can admit a non-local UV completion only if the scale of non-locality $M$ is of the same order or lower than the EFT cutoff.
  • On the other hand, our positivity bounds still allow for large values of the scale $\alpha$ (i.e. $\alpha\Lambda^{4}>1.02$). However, these values might not be compatible with partial-wave unitarity of the full amplitude. Although we do not have a proof, some examples supporting this statement are provided in App. A, where we notice that partial-wave unitarity is satisfied only if the EFT cutoff is close to $\alpha^{-1/4}$. Insights from IR causality might suggest this situation is generic.
Finally, it is worth mentioning that non-local UV completions whose amplitudes are compatible with the properties of exponential boundedness, unitarity, analyticity and IR causality, may still show some causality violation at high energies. Our study cannot say anything about UV causality, but this is an aspect that will be investigated in future work.
6 Conclusions
In this paper, we derived generalized positivity bounds on the Wilson coefficients of the EFT expansion assuming that the low-energy theory admits a non-local UV completion. In particular, we focused on a scalar theory. The central object of our discussion was a $2\rightarrow 2$ elastic scattering amplitude satisfying the properties of analyticity, crossing symmetry, and partial-wave unitarity, but not polynomial boundedness. Indeed, we considered amplitudes that are exponentially bounded in the complex energy plane, which are typical of non-local QFTs.
To compensate for the exponential growth of the amplitude we derived modified dispersion relations. While this was straightforward in the forward limit, for non-zero momentum transfer we had to require a specific but still quite generic crossing-symmetric form of the amplitude. By doing this, we successfully ensured that the integral over the arc at infinity did not diverge but vanished, thus suitable dispersion relations could be used. Consequently, we derived “non-local positivity bounds” and compared them with the standard local ones. We showed that the allowed regions in the space of EFT parameters can either overlap or not with those in the local case, depending on the value of the ratio of the non-locality scale to EFT cutoff. In particular, we found regions where the corresponding EFTs do not admit a local UV completion, but only a non-local one.
We also compared the new positivity bounds with low-energy causality constraints [71]. We found that causal EFTs can admit non-local UV completions only if the scale of non-locality – defined as $M=\alpha^{-1/4}$ – is of the same order or smaller than the EFT cutoff $\Lambda$, i.e. if the following inequality is satisfied: $M<(1.02)^{1/4}\Lambda.$
To support our main findings, in app. A we also presented an explicit example that serves as a proof of concept for the existence of a non-perturbative amplitude that satisfies all the standard desired properties (i.e. partial-wave unitarity, causality, crossing symmetry) except for the polynomial boundedness.
It is also worth to mention that non-local (or non-localizable) theories can be relevant for the construction of a UV completion of general relativity in the form of four-dimensional QFTs, such as infinite derivative gravity [91], or asymptotically safe gravity formulated in terms of form factors in the quantum effective action [92, 93]. Indeed, it would also be interesting to compare our positivity bounds with constraints derived in the context of asymptotically safe theories [94, 95, 96, 97]. Our work reinforce the idea that these quantum-gravity approaches could be consistent with the full set of unitarity constraints and with macroscopic causality. Exploring and classifying different types of viable exponentially bounded amplitudes could be a very interesting direction for future studies. This has the potential to shed new light on the form of the quantum effective action in UV-complete gravitational theories, as well as on the question of the uniqueness of string theory as the only consistent UV completion of general relativity.
Acknowledgements
Authors are indebted to Ivano Basile, Gia Dvali, Alessia Platania, Xinan Zhou, and Shuang-Yong Zhou for several enlightening conversations. A. T. thanks Alexander Zhiboedov for his hospitality at CERN and numerous deep discussions on the reconstruction of the unitary and causal amplitudes. A. T. thanks GuangZhuo Peng and YoungJun Xu for useful comments and feedback. The authors are grateful to Marianna Carillo-Gonzales for sharing her numerical results on IR causality bound. L. B. and A. T. thank The Royal Society for financial support through the International Exchanges Scheme 2022. L. B. acknowledges financial support from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Actions (grant agreement ID: 101106345-NLQG). The work of A. T. was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 1234710.
Appendix A Example of unitary and causal amplitude violating Froissart-Martin bound
Scattering amplitudes that are exponentially bounded in the complex energy plane can be obtained, for example, in theories containing infinite derivative operators [98]. This type of non-polynomial behavior is usually associated to the lack of locality of the underlying field theory. It is still under debate whether these theories are consistent with the other standard fundamental requirements, such as causality. Here we construct working examples of UV complete amplitudes that are exponentially bounded but still satisfy partial-wave unitarity, crossing symmetry and causality.
The simplest form of such an amplitude could be
$A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}\,.$ (35)
However, it can be shown that this expression does not satisfy partial-wave unitarity for large $s$ due to the fact that the corresponding partial waves heavily oscillate and can give rise to negative imaginary contributions. In other words, if we compute
${\rm Disc}A(s,t,u)\propto e^{-\alpha s^{2}}+e^{-\alpha t^{2}}+e^{-\alpha u^{2}}$ (36)
and expand in partial waves, it can be shown that the imaginary part of the partial waves can be negative for large values of $s.$
However, we can try to suitably modify the amplitude and check whether we can find expressions that are compatible with unitarity. For example, it can be shown that the following discontinuity,
${\rm Disc}_{s}A(s,t,u)\propto-\left(e^{-\alpha s^{2}}\log{s}+e^{-\alpha t^{2}}\log{t}+e^{-\alpha u^{2}}\log{u}\right)\,,$ (37)
has positive-definite $s$-discontinuities of the partial waves. A possible form of the amplitude whose discontinuity is of the type in eq. (37) but is defined in the physical region (i.e. $t\leq 0$) is given by
$A(s,t,u)=$ $C\left[e^{-\alpha s^{2}}\log{s}\log{(-s)}+e^{-\alpha t^{2}}\log{t}\log{(-t)}+e^{-\alpha u^{2}}\log{u}\log{(-u)}\right]$ (38)
$=$ $C\left[e^{-\alpha s^{2}}\log{(s)}\,(\log{s}-i\pi)+e^{-\alpha t^{2}}\log{(-t)}\,(\log{(-t)}-i\pi)+\right.$
$\qquad\qquad\qquad\left.+e^{-\alpha u^{2}}\,\log{(-u)}(\log{(-u)}-i\pi)\right]\,.$
Here the branch cut of the logarithm is specified by the condition $\log{s}-\log{(-s)}=i\pi$. This amplitude can be shown to satisfy partial-wave unitary for reasonable values of the pre-factor $C.$ The behavior of the real and imaginary parts of some partial waves are plotted in Fig. 4a for $C=1$.
A weak point the amplitude (38) is that some of the partial waves blow up in the $s\rightarrow 0$ limit (see Fig. 4a). We can still find a working example with all imaginary parts of the partial waves that are positive and finite. This is given by the amplitude
$A(s,t,u)=C\left(s^{\gamma i}e^{-\alpha s^{2}}+t^{\gamma i}e^{-\alpha t^{2}}+u^{\gamma i}e^{-\alpha u^{2}}\right)\,.$ (39)
If $\gamma$ is small, the power can be expanded in terms of logarithms, so it will behave similarly to (38). In what follows we show that for this amplitude unitarity and causality requirements are satisfied for all real positive $s$. Therefore, this can serve as a proof of concept for the existence of a consistent exponentially bounded amplitude that is compatible with the crossing symmetric structure in (18), which was used in this paper to derive positivity bounds.
Partial-wave unitarity check.
The unitarity condition on the scattering amplitude formulated in terms of the partial waves reads
$2\,{\rm Im}f_{l}\left(s\right)\geq\left|f_{l}\left(s\right)\right|^{2}\,,$ (40)
where in four spacetime dimensions $f_{l}(s)$ can be computed as
$f_{l}\left(s\right)=\dfrac{1}{16\pi}\int^{1}_{-1}{\rm d}xP_{l}\left(x\right)A\left(s,-\dfrac{s}{2}\left(1-x\right),-\dfrac{s}{2}\left(1+x\right)\right).$ (41)
Figure 4: (a) Real (solid lines) and imaginary (dashed lines) parts of the partial-wave amplitudes as functions of $s$ for the amplitude (38). (b) Real (solid lines) and imaginary (dashed lines) parts of the partial-wave amplitudes as functions of $s$ for the amplitude (39) with $\gamma=0.1$. We set $C=1$ in both figures.
For the amplitude (39), we can write an analytic series expression of the partial waves for $l>0$ by expanding the exponent and integrating term by term, i.e.
$f_{l}(s)=\dfrac{1}{4\pi}\sum^{\infty}_{n=0}\frac{\left(-\alpha\right)^{n}\left(-s\right)^{2ng+i\gamma}}{\Gamma(n+1)}\dfrac{\left(\Gamma\left(i\gamma+2ng+1\right)\right)^{2}}{\Gamma\left(i\gamma+2ng+l+2\right)\Gamma\left(i\gamma+2ng-l+1\right)}.$ (42)
From this expression it is evident that the partial waves have the same type of essential singularity at infinity of the full amplitude, that is, they are not polynomially bounded in a complex energy plane. However, the unitarity condition is required only for physical momenta $s\geq 0$, which means that their growth in the complex energy plane is not necessarily in contradiction with partial-wave unitarity.
The sum in the Eq. (42) can be expressed in terms of rather complicated hypergeometric functions, but we noticed that it was more efficient to check unitarity by computing partial waves numerically from (41). We found that the amplitude (39) satisfies the positivity requirement ${\rm Im}f_{l}(s)>0$ if $\gamma\leq 0.1$. As all partial waves are bounded, the full non-linear unitarity condition can be always achieved by suitably adjusting the overall constant factor. In Fig. 4b we plotted the numerical solutions for $\gamma=0.1.$
A probe of causality.
The notion of causality is typically encoded in an analyticity requirement for the scattering amplitudes [56]. However, analyticity itself does not guarantee the absence of superluminal propagation. The condition forbidding any time advance in the propagation of a signal (also called asymptotic causality) requires the following inequality to be valid [99, 100]:
$\Delta T=2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}\,\delta_{l}(s)\approx 2\sqrt{s}\frac{\partial}{\partial\sqrt{s}}\text{Re}f_{l}(s)>0\,,\qquad f_{l}(s)=-i(e^{i\delta_{l}(s)}-1).$ (43)
Here $\delta_{l}$ is a scattering phase corresponding to the partial wave amplitude, and the approximation used means that $f_{l}(s)$ is small.
Due to oscillations in the real part of the partial waves for small values of $s$ (see Fig. 4b), this strict causality condition cannot be satisfied. However, the condition of macroscopic causality that still allows for unresolvable time advances (33) can instead be respected. This condition is more justified from a quantum mechanical point of view and is given by
$\sqrt{s}\Delta T(s)=4s\frac{\partial}{\partial s}\text{Re}f_{l}(s)>-1.$ (44)
In Fig. 5 we showed the behavior of $\sqrt{s}\Delta T(s)$ for the amplitude in (39). We found that indeed macroscopic causality is satisfied everywhere from zero to infinite energies since $|\sqrt{s}\Delta T(s)|<1$ for any $s$.
Figure 5: Plot of $\sqrt{s}\Delta T$ for the amplitude (39) with $\gamma=0.1$. This quantity is bounded from below and from above, and never becomes smaller than $-1$. This is consistent with the causality condition that forbids resolvable time advances. However, in this case, asymptotic causality is not satisfied.
Appendix B Derivation of the analytic bound on $g_{3}$
In this appendix, we provide more details about the method that was used to derive the bounds involving the coupling $g_{3}$. We refer to [101, 69, 102] for more detailed derivation and explanations.
If we define
$a_{k,q}=\left[\frac{L^{2q}}{\mu^{k+1}}\right]=\sum_{l\,\rm{even}}n_{l}\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\rho_{l}(\mu)e^{\alpha\mu^{2}}\frac{L^{2q}}{\mu^{k+1}}\,,$ (45)
we can write the following relaxing equalities:
$a_{2,0}-a_{3,0}\geq 0\,,\quad a_{3,0}-a_{4,0}\geq 0\,,\quad a_{3,1}-a_{4,1}\geq 0\,,\dots\,.$ (46)
Calling $f_{1}(\mu,L^{2})=1/\mu^{3/2},\,f_{2}(\mu,L^{2})=1/\mu^{5/2},\,f_{3}(\mu,L^{2})=L^{2}/\mu^{5/2},$ and applying the Gram’s inequality [103], we obtain
$\left|\begin{matrix}\int f_{1}^{2}&\int f_{1}f_{2}&\int f_{1}f_{3}\\[4.2679pt] \int f_{2}f_{1}&\int f_{2}^{2}&\int f_{2}f_{3}\\[4.2679pt] \int f_{3}f_{1}&\int f_{3}f_{2}&\int f_{3}^{2}\end{matrix}\right|\geq 0\quad\Rightarrow\quad\left|\begin{matrix}a_{2,0}&a_{3,0}&a_{3,1}\\[4.2679pt] a_{3,0}&a_{4,0}&a_{4,1}\\[4.2679pt] a_{3,1}&a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,,$ (47)
where for brevity we omitted the measure ${\rm d}\mu$ and the integration range $[\Lambda^{2},\infty]$ in the above integrals. In addition, we also have the rank-2 version of Gram’s inequalities,
$\left|\begin{matrix}a_{2,0}&a_{3,0}\\[4.2679pt] a_{3,0}&a_{4,0}\\ \end{matrix}\right|\geq 0\,,\qquad\left|\begin{matrix}a_{2,0}&a_{3,1}\\[4.2679pt] a_{3,1}&a_{4,2}\end{matrix}\right|\geq 0\,,\qquad\left|\begin{matrix}a_{4,0}&a_{4,1}\\[4.2679pt] a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,.$ (48)
The final condition follows from the fact that $L^{2}\in\{0,6,20,42,\dots\}$ is discrete-valued, thus we have $(L^{2}-6)(L^{2}-20)\geq 0$ for all possible values of $L^{2}$ in the discrete set. Thus, we have
$\left[\frac{(L^{2}-6)(L^{2}-20)}{\mu^{5}}\right]\geq 0,\qquad a_{4,2}-26a_{4,1}+120a_{4,0}\geq 0.$ (49)
To obtain the constrained regions in the $a_{3,1}\text{-}a_{4,0}$ plane (corresponding to $g_{3}\text{-}g_{4}$ plane), we need to project out $a_{4,2},$ $a_{4,1}$ and $a_{3,0}$. The null constraint (26) can be written as $a_{4,2}=8a_{4,1}$ and can help us to eliminate $a_{4,2}$. Moreover, to project out $a_{4,1}$ and $a_{3,0}$ it is essential to saturate the most “loose” bounds containing $a_{4,1}$. For example, after using the null constraint, Eq. (49) becomes
$a_{4,0}\leq\frac{20}{3}a_{4,1}\,.$ (50)
Thus, we take $a_{4,0}=20a_{4,1}/3$ to eliminate $a_{4,1}$. Consequently, Eq. (47) gives
$\tilde{a}_{3,1}^{2}-\frac{40}{3}\tilde{a}_{3,0}\tilde{a}_{3,1}+\frac{160}{3}\tilde{a}_{3,0}^{2}-\frac{80}{9}\tilde{a}_{4,0}\leq 0\,,$ (51)
where $\tilde{a}_{k,q}\equiv\frac{a_{k,q}}{a_{2,0}}\Lambda^{2(k-2)}$. The last inequality implies
$\tilde{a}_{3,1}\leq\frac{20}{3}\tilde{a}_{3,0}+\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{3,0}^{2})}\,.$ (52)
If we define the function $f(\tilde{a}_{3,0})\equiv\frac{20}{3}\tilde{a}_{3,0}+\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{3,0}^{2})}$, then the most loose bound corresponds the maximum of $f(\tilde{a}_{3,0})$. This maximum value is given by $\tilde{a}_{3,0}=\sqrt{\frac{5}{6}\tilde{a}_{4,0}}$. In addition, from the other constraints in Eqs. (46) and (48), we also have $\tilde{a}_{4,0}\leq\tilde{a}_{3,0}\leq\sqrt{\tilde{a}_{4,0}}$, which means that $f(\tilde{a}_{3,0})$ can also reach maximum at the boundary of the allowed range for $\tilde{a}_{3,0}$. Therefore, we need to consider these two regions separately,
$0<\tilde{a}_{3,1}<\sqrt{\frac{160}{3}\tilde{a}_{4,0}}$ $\text{for}\qquad 0<\tilde{a}_{4,0}\leq\frac{5}{6},$ (53)
$0<\tilde{a}_{3,1}<\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{4,0}^{2})}+\frac{20}{3}\tilde{a}_{4,0}$ $\text{for}\qquad\frac{5}{6}<\tilde{a}_{4,0}\leq 1.$
The shape carved by these inequalities is called a-geometry, the result is general, and does not rely on the definition of the brackets in Eq. (45). In our study that involves non-local amplitudes, we need to replace $\tilde{a}_{4,0}=2\tilde{g}_{4}+\tilde{\alpha}$ and $\tilde{a}_{3,1}=-\frac{1}{2}\tilde{g}_{3}+\frac{3}{2}\tilde{a}_{3,0}$ into Eq. (51) and then project out $\tilde{a}_{3,0}$ to get the result in Eq. (29).